OpenStudy (anonymous):

Let g(x)=x^2+2x-6. Find 2[g(2)] I'm lost, how would I do this?? (Sorry, my brain isn't working tonight...)

OpenStudy (zale101):

Find 2[g(2)] 2 is multiplied by g(2) What's g(2)? It means the function g(x)=x^2+2x-6 will have an input of 2 instead of x so solve this g(2)= then multiply the results by 2

OpenStudy (zale101):

g(2)=(2)^2+2(2)-6 =?

OpenStudy (anonymous):

g(2)=2?

OpenStudy (zale101):

yes, so we know that g(2)=2 2[g(2)] 2[2]=4

OpenStudy (anonymous):

Thank you again! :)

OpenStudy (zale101):

You're Welcome~